# Debiased Machine Learning and Network Cohesion for Doubly-Robust Differential Reward Models in Contextual Bandits

### v2

@@ E.1 Data from the Valentine Study
Feature Description Interaction Baseline Model
@@
Prior week step count variability SD of the rewards in prior week $\times$ $\surd$
@@ E.2 Evaluation
We generated 100 bootstrap samples and ran our contextual bandit on them, considering the binary action of whether or not to send a message at a given decision point based on the contextual variables $S_{1}$ and $S_{2}$. Each user is considered independently and with a cohesion network, for maximum personalization and independence of results. To guarantee that messages have a positive probability of being sent, we only sample the observations with notification randomization probability between $0.01$ and $0.99$. In the case of the algorithm utilizing NNR, we chose four baseline characteristics (gender, age, device, and baseline average daily steps) to establish a measure of “distance” between users. For this analysis, the value of $k$ representing the number of nearest neighbors was set to 5. To utilize bootstrap sampling, we train the Neural-Linear method’s neural network using out-of-bag samples. The neural network architecture comprises a single hidden layer with two hidden nodes. The input contains both the baseline characteristics and the contextual variables and the activation function applied here is the softplus function, defined as $\text{softplus}(x)=\log(1+\exp{(x)})$.
@@
E.3 Inverse Propensity Score (IPS) offline evaluation
@@
Appendix F Additional Details for the Intern Health Study (IHS)
To further enhance the competitive performance of our proposed DML-TS-NNR algorithm, we conducted an additional comparative analysis using a real-world dataset from the Intern Health Study (IHS) (NeCamp et al., 2020). This micro-randomized trial investigated the use of mHealth interventions aimed at improving the behavior and mental health of individuals in stressful work environments. The estimates obtained represent the improvement in average reward relative to the original constant randomization, averaging across stages (K = 30) and participants (N = 1553). The available IHS data consist of 20 multiple-imputed data sets. We apply the algorithms to each imputed data set and perform a comparative analysis of the competing algorithms. The results presented in Figure 7 shows our proposed DML-TS-NNR-RF algorithm achieved significantly higher rewards than the other three competing ones and demonstrated comparable performance to the AC algorithm. These findings further support the advantages of our proposed algorithm.
F.1 Data from the IHS
@@
Feature
Description
Interaction Baseline
Day in study
an integer from $1$ to $30$
$\surd$ $\surd$
Average daily steps in prior five days
cubic root
$\surd$ $\surd$
Average daily sleep in prior five days
cubic root
$\times$ $\surd$
Average daily mood in prior five days
a Likert scale from $1-10$
$\times$ $\surd$
Pre-intern average daily steps
cubic root
$\times$ $\surd$
Pre-intern average daily sleep
cubic root
$\times$ $\surd$
Pre-intern average daily mood
a Likert scale from $1-10$
$\times$ $\surd$
Sex
Gender
$\times$ $\surd$
Week category
The theme of messages in a specific week (mood, sleep, activity, or none)
$\times$ $\surd$
PHQ score
PHQ total score
$\times$ $\surd$
Early family environment
higher score indicates higher level of adverse experience
$\times$ $\surd$
Personal history of depression
$\times$ $\surd$
Neuroticism (Emotional experience)
higher score indicates higher level of neuroticism
$\times$ $\surd$
@@
At every decision point, before selecting an action, the learner sees two state variables: the participant’s previous 5-day average daily step count ($S_{1}$, cubic root) and the participant’s day in study ($S_{2}$, an integer from $1$ to $30$).
F.2 Evaluation
@@
We run our contextual bandit on the IHS data, considering the binary action of whether or not to send a message at a given decision point based on the contextual variables $S_{1}$ and $S_{2}$. Each user is considered independently and with a cohesion network, for maximum personalization and independence of results. To guarantee that messages have a positive probability of being sent, we only sample the observations with notification randomization probability between $0.01$ and $0.99$. For the algorithm employing NNR, we defined participants in the same institution as their own “neighbors”. This definition enables the flexibility for the value of $k$, representing the number of nearest neighbors, to vary for each participant based on their specific institutional context. Furthermore, in our study setting, we make the assumption that individuals from the same ”institution” enter the study simultaneously as a group. Due to the limited access to prior data, we are unable to build the neural linear models as in the Valentine Study.
We utilized 20 multiple-imputed data sets and performed an offline evaluation of the contextual bandit algorithms on each data set. The result is presented below in Figure 7.
@@
Figure 7: (left) Unbiased estimates of the average per-trial reward for all five competing algorithms, relative to the reward obtained under the pre-specified Valentine randomization policy across 20 multiple-imputed data sets. And (right) p-values from the pairwise paired t-tests. The dark shade in the last column indicates that the proposed DML-TS-NNR-RF algorithm achieved significantly higher rewards than the other three competing algorithms while demonstrating comparable performance to the AC algorithm.
Appendix G Regret Bound
G.1 Double Robustness of Pseudo-Reward
@@ G.2 Preliminaries
so that plugging this into Eqn. 9, we have
@@
and plugging this into Eqn. 8 we obtain the variance.
@@
and plugging this into Eqn. 10,
@@
[…]
@@
[…]
@@
[…]
We define denote the rhs as $v_{k}^{2}$ e.g.
[…]
@@
$C$ $\equiv\max\left(A,16B^{2}\right).$
@@
[…]
@@
Here we collect three important results. For analysis at stage $K$, let $\phi(x_{i,t})$ encodes the vector $x_{i,t}$ for the $i$th individual and the $t$th time in a vector of length $2d*K$. At stage $k$, let $\mathcal{O}_{k}$ denote the set of observed time points across all individuals at stage $k$. Let $\hat{\theta}_{k}$ denote the estimates for stage $k+1$ using data from $\mathcal{O}_{k}$. First, we prove a slightly modified version of Lemma 10 from Abbasi-Yadkori et al. (2011):
Lemma 6.
(Determinant-Trace Inequality) Suppose $\tilde{\sigma}_{i,t}\phi(\boldsymbol{x}_{i,t})\in\mathbb{R}^{d}$. Let $V_{k+1}=\sum_{(i,t)\in\mathcal{O}_{k}}\tilde{\sigma}_{i,t}^{2}\phi(\boldsymbol{x}_{i,t})\phi(\boldsymbol{x}_{i,t})^{\top}+V_{0}$. Then
[…]
Proof.
Following the arguments for the proof of the original Lemma 11 in Abbasi-Yadkori et al. (2011), we have
$\textrm{det}(V_{k+1})$ $\leq\left(\frac{\textrm{tr}(V_{k+1})}{d}\right)^{d}.$
Now we have $\textrm{tr}(\tilde{\sigma}_{i,t}^{2}\phi(\boldsymbol{x}_{i,t})\phi(\boldsymbol{x}_{i,t})^{\top})\leq 1/4$ since $\|x_{i,t}\|\leq 1$ and $\tilde{\sigma}_{i,t}^{2}\leq(1/2)^{2}$.
$\textrm{tr}(V_{k+1})$ $=\textrm{tr}(V_{0})+\sum_{(i,t)\in\mathcal{O}_{k}}\tilde{\sigma}_{i,t}^{2}\textrm{tr}(\phi(\boldsymbol{x}_{i,t})\phi(\boldsymbol{x}_{i,t})^{\top})$
$=\textrm{tr}(V_{0})+\sum_{(i,t)\in\mathcal{O}_{k}}\tilde{\sigma}_{i,t}^{2}\|\phi(\boldsymbol{x}_{i,t})\|^{2}_{2}$
$\leq\textrm{tr}(V_{0})+\frac{1}{4}|\mathcal{O}_{k}|$
Plugging this in to the rhs of the determinant inequality in the first step gives the desired result. ∎
We adapt an important concentration inequality for regularized least-squares estimates. Our proof follows the same basic strategy as Abbasi-Yadkori et al. (2011), but with modifications due to 1) the use of weighted least squares 2) the use of pseudo-rewards to estimate differential rewards 3) replacing the scaled diagonal regularization with Laplacian regularization.
@@
For any $\delta>0$, w.p. at least $1-\delta$ the estimates $\{\hat{\Theta}_{k}\}_{k=0}^{\infty}$ in Algorithm 1 satisfies for any $\{x_{k}\}_{k=0}^{\infty}$,
$|x_{k}^{\top}(\hat{\Theta}_{k}-\Theta^{*}_{k})|$ $\leq\|x_{k}\|_{V_{k-1}^{-1}}\left(v_{k}\sqrt{2\log\left(\frac{\det(V_{k-1})^{1/2}\det(V_{0})^{-1/2}}{\delta}\right)}+\lambda_{\max}(V_{0})^{1/2}kB\right),$ (11)
where $v_{k}^{2}$ is the variance factor for the difference between the pseudo-reward and its mean at stage $k$. In particular, setting $x_{k}=V_{k-1}(\hat{\Theta}_{k-1}-\Theta_{k}^{*})$ implies
$\|\hat{\Theta}_{k}-\Theta_{k}^{*}\|$ $\leq v_{k}\sqrt{2\log\left(\frac{\det(V_{k-1})^{1/2}\det(V_{0})^{-1/2}}{\delta}\right)}+\lambda_{\max}(V_{0})^{1/2}kB$
@@
$=V_{k-1}^{-1}\xi_{k}+V_{k-1}^{-1}\Phi_{k}^{\top}W_{k}\Phi_{k}\Theta_{k}^{*}$
$=V_{k-1}^{-1}\xi_{k}+V_{k-1}^{-1}(\Phi_{k}^{\top}W_{k}\Phi_{k}+V_{0})\Theta_{k}^{*}-V_{k-1}^{-1}V_{0}\Theta_{k}^{*}$
$=V_{k-1}^{-1}\xi_{k}+\Theta_{k}^{*}-V_{k-1}^{-1}V_{0}\Theta_{k}^{*}$
@@
$\hat{\Theta}_{k}-\Theta^{*}_{k}$ $=V_{k-1}^{-1}(\xi_{k}-V_{0}\Theta_{k}^{*})$
@@
$|x_{k}^{\top}\hat{\Theta}_{k}-x_{k}^{\top}\Theta_{k}^{*}|$ $\leq\|x_{k}\|_{V_{k-1}^{-1}}(\|\xi_{k}\|_{V_{k-1}^{-1}}+\|V_{0}\Theta_{k}^{*}\|_{V_{k-1}^{-1}})$
@@
$\|V_{0}\Theta_{k}^{*}\|_{V_{k-1}^{-1}}^{2}$ $=\Theta_{k}^{*\top}V_{0}^{\top}V_{k-1}^{-1}V_{0}\Theta_{k}^{*}$
$\leq\|V_{0}^{\top}V_{k-1}^{-1}V_{0}\|_{2}\|\Theta_{k}^{*}\|_{2}^{2}$
$\leq\|V_{0}\|^{2}_{2}\|V_{k-1}^{-1}\|_{2}\|\Theta_{k}^{*}\|_{2}^{2}$
$\leq\lambda_{\textrm{max}}(V_{0})\|\Theta_{k}^{*}\|_{2}^{2}$
@@
$|x_{k}^{\top}\hat{\Theta}_{k}-x_{k}^{\top}\Theta_{k}^{*}|$ $\leq\|x_{k}\|_{V_{k-1}^{-1}}\left(v_{k}\sqrt{2\log\left(\frac{\textrm{det}(V_{k-1}^{-1})^{1/2}\textrm{det}(V_{0})^{-1/2}}{\delta}\right)}+\lambda_{\max}(V_{0})^{1/2}kB\right)$
@@
$|x_{k}^{\top}(\hat{\Theta}_{k}-\Theta^{*}_{k})|$ $\leq\|x_{k}\|_{V_{k-1}^{-1}}\left(\left(\frac{C}{k^{1/2}}+\sigma^{2}c^{2}\right)\sqrt{2\log\left(\frac{\det(V_{k-1})^{1/2}\det(V_{0})^{-1/2}}{\delta/2}\right)}+\lambda_{\max}(V_{0})^{1/2}kB\right),$ (12)
In particular, setting $x_{k}=V_{k-1}(\hat{\Theta}_{k-1}-\Theta_{k}^{*})$ implies
$\|\hat{\Theta}_{k}-\Theta_{k}^{*}\|$ $\leq\left(\frac{C}{k^{1/2}}+\sigma^{2}c^{2}\right)\sqrt{2\log\left(\frac{\det(V_{k-1})^{1/2}\det(V_{0})^{-1/2}}{\delta/2}\right)}+\lambda_{\max}(V_{0})^{1/2}kB$
@@
Let $\lambda\geq 1$ and $\gamma\geq 1$. For any arbitrary sequence $(\boldsymbol{x}_{i,t})_{(i,t)\in\mathcal{O}_{k}}$, let
$V_{k+1}\equiv\sum_{(i,t)\in\mathcal{O}_{k}}\tilde{\sigma}_{i,t}^{2}\phi(\boldsymbol{x}_{i,t})\phi(\boldsymbol{x}_{i,t})^{\top}+V_{0},$
@@
$\sum_{k=1}^{K}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\|\phi(\boldsymbol{x}_{i,t})\|_{V_{k}^{-1}}^{2}\leq 2c\log\left(\frac{\det(V_{K+1})}{\det(V_{0})}\right).$
@@
$\sum_{k=1}^{K}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\tilde{\sigma}_{i,t}^{2}\|\phi(\boldsymbol{x}_{i,t})\|_{V_{k}^{-1}}^{2}\leq 2\log\left(\frac{\det(V_{K+1})}{\det(V_{0})}\right).$
@@
$\sum_{k=1}^{K}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\|\phi(\boldsymbol{x}_{i,t})\|_{V_{k}^{-1}}^{2}$ $\leq 2c\log\left(\frac{\det(V_{K+1})}{\det(V_{0})}\right)$
as desired.
By definition, we have $tr(\tilde{\sigma}_{i,t}^{2}\phi(\boldsymbol{x}_{i,t})\phi(\boldsymbol{x}_{i,t})^{\top})\leq 2/4$ since $\|x_{i,t}\|\leq 1$ and $\tilde{\sigma}_{i,t}^{2}\leq(1/2)^{2}$; and $tr(L_{\otimes})=tr(L)\cdot tr(I)=KMd$. Then by Hadamard’s inequality we have
@@
$=2Kd\log\left(\frac{1}{2Kd}\left[\gamma 2Kd+\lambda 2KMd+\frac{2}{4}K\cdot(K+1)/2\right]\right)$
$=2Kd\log\left(\left[\gamma+\lambda M+\frac{K+1}{8d}\right]\right)$

### v3

@@ C.1 Data from the Valentine Study
Feature Description Interaction Baseline
@@
Prior week step count variability SD of the rewards in the previous week $\times$ $\surd$
@@ C.2 Evaluation
We generated 100 bootstrap samples and ran our contextual bandit on them, considering the binary action of whether or not to send a message at a given decision point based on the contextual variables $S_{1}$ and $S_{2}$. Each user is considered independently and with a cohesion network, for maximum personalization and independence of results. To guarantee that messages have a positive probability of being sent, we only sample the observations with notification randomization probability between $0.01$ and $0.99$. In the case of the algorithm utilizing NNR, we chose four baseline characteristics (gender, age, device, and baseline average daily steps) to establish a measure of “distance" between users. For this analysis, the value of $k$ representing the number of nearest neighbors was set to 5. To utilize bootstrap sampling, we train the Neural-Linear method’s neural network using out-of-bag samples. The neural network architecture comprises a single hidden layer with two hidden nodes. The input contains both the baseline characteristics and the contextual variables and the activation function applied here is the softplus function, defined as $\text{softplus}(x)=\log(1+\exp{(x)})$.
@@
Figure 9: (left) Unbiased estimates of the average per-trial reward for all three ablation algorithms, relative to the reward obtained under the pre-specified Valentine Study randomization policy across 20 multiple-imputed data sets. And (right) p-values from the pairwise paired t-tests.
C.3 An ablation study
We further investigate the primary contributors to the algorithm’s performance, we conducted a parallel ablation study using real-world data analysis. In this study, we compared DML-TS-NNR (as detailed in the main paper) with two additional methods for reference: DML-TS-SU and NNR-Linear. Each comparison focuses on specific elements of our algorithm to discern their impact on DML-TS-NNR’s overall performance.
The differences between these methods are illustrated in Figure. 9. In this particular dataset, NNR plays a more significant role in driving the overall superior performance of our algorithm.
C.4 Inverse Propensity Score (IPS) offline evaluation
@@
Appendix D Additional Details for the Intern Health Study (IHS)
To further enhance the competitive performance of our proposed DML-TS-NNR algorithm, we conducted an additional comparative analysis using a real-world dataset from the Intern Health Study (IHS) (NeCamp et al., 2020). This micro-randomized trial investigated the use of mHealth interventions aimed at improving the behavior and mental health of individuals in stressful work environments. The estimates obtained represent the improvement in average reward relative to the original constant randomization, averaging across stages (K = 30) and participants (N = 1553). The available IHS data consist of 20 multiple-imputed data sets. We apply the algorithms to each imputed data set and perform a comparative analysis of the competing algorithms. The results presented in Figure 10 shows our proposed DML-TS-NNR algorithm achieved significantly higher rewards than the other three competing ones and demonstrated comparable performance to the AC algorithm. These findings further support the advantages of our proposed algorithm.
D.1 Data from the IHS
@@
Feature Description Interaction Baseline
Day in study an integer from $1$ to $30$ $\surd$ $\surd$
Average daily steps in prior five days cubic root $\surd$ $\surd$
Average daily sleep in prior five days cubic root $\times$ $\surd$
Average daily mood in prior five days a Likert scale from $1-10$$\times$ $\surd$
Pre-intern average daily steps cubic root $\times$ $\surd$
Pre-intern average daily sleep cubic root $\times$ $\surd$
Pre-intern average daily mood a Likert scale from $1-10$$\times$ $\surd$
Sex Gender $\times$ $\surd$
Week category The theme of messages in a specific week (mood, sleep, activity, or none) $\times$ $\surd$
PHQ score PHQ total score $\times$ $\surd$
Early family environment higher score indicates higher level of adverse experience $\times$ $\surd$
Personal history of depression $\times$ $\surd$
Neuroticism (Emotional experience) higher score indicates higher level of neuroticism $\times$ $\surd$
@@
At every decision point, before selecting an action, the learner sees two state variables: the participant’s previous 5-day average daily step count ($S_{1}$, cubic root) and the participant’s day in the study ($S_{2}$, an integer from $1$ to $30$).
D.2 Evaluation
@@
We run our contextual bandit on the IHS data, considering the binary action of whether or not to send a message at a given decision point based on the contextual variables $S_{1}$ and $S_{2}$. Each user is considered independently and with a cohesion network, for maximum personalization and independence of results. To guarantee that messages have a positive probability of being sent, we only sample the observations with notification randomization probability between $0.01$ and $0.99$. For the algorithm employing NNR, we defined participants in the same institution as their own “neighbors". This definition enables the flexibility for the value of $k$, representing the number of nearest neighbors, to vary for each participant based on their specific institutional context. Furthermore, in our study setting, we assume that individuals from the same "institution" enter the study simultaneously as a group. Due to the limited access to prior data, we are unable to build the neural linear models as in the Valentine Study.
We utilized 20 multiple-imputed data sets and performed an offline evaluation of the contextual bandit algorithms on each data set. The result is presented below in Figure 10. Similar to Section C.3, here we also compared DML-TS-NNR (as detailed in the main paper) with DML-TS-SU and NNR-Linear. The differences between these methods are illustrated in Figure. 11. In this specific dataset, both NNR and DML exhibit comparable performance individually, but their combined effect significantly enhances the overall performance of the algorithm.
@@
Figure 10: (left) Unbiased estimates of the average per-trial reward for all four competing algorithms, relative to the reward obtained under the pre-specified Intern Health Study randomization policy across 20 multiple-imputed data sets. And (right) p-values from the pairwise paired t-tests. The dark shade in the last column indicates that the proposed DML-TS-NNR algorithm achieved significantly higher rewards than the other three competing algorithms while demonstrating comparable performance to the AC algorithm.
Figure 11: (left) Unbiased estimates of the average per-trial reward for all three ablation algorithms, relative to the reward obtained under the pre-specified Intern Health Study randomization policy across 20 multiple-imputed data sets. And (right) p-values from the pairwise paired t-tests.
Appendix E Additional Details for Algorithm 1
This appendix briefly discusses two details of Algorithm 1. The first is efficient computation of $V^{-1}$. Because $V$ is a fairly large matrix, a full matrix inversion is expensive. Fortunately, however, we can dramatically reduce the necessary computational requirements because each additional $(i,t)$ involves a rank-one perturbation to $V$. Consequently, we can apply the Sherman–Morrison formula to speed up computations. We leveraged this trick in our implementation of DML-TS-NNR for the simulation study by using efficient rank-one updates available in the SuiteSparse library (Davis & Hu, 2011).
The second detail is the requirements for the distribution, $\mathcal{D}^{TS}$:
Definition 1.
While a Gaussian prior satisfies Definition 1, this approach allows us to move beyond Bayesian posteriors to generic randomized policies. As discussed in the main paper, we recommend setting $\mathcal{D}^{TS}$ to a multivariate Gaussian distribution or multivariate t-distribution. These choices simplify computations because the probability computation in (10) simplifies to an evaluation of the CDF of either (a) a univariate Gaussian or (b) a univariate t-distribution. The mean and variance of the corresponding univariate distribution can easily be worked out using the moments of a linear combination of random variables.
Appendix F Regret Bound
F.1 Double Robustness of Pseudo-Reward
@@ F.2 Preliminaries
so that plugging this into Eqn. 14, we have
@@
and plugging this into Eqn. 13 we obtain the variance.
@@
and plugging this into Eqn. 15,
@@
$\mathbb{E}_{p(s,\bar{a})}\left[\textrm{Var}\left(\tilde{R}_{i,t}^{f}(s,\bar{a})\right)\right]$ $\leq\frac{\mathbb{E}_{p(s,\bar{a})}\left[(r_{i,t}(s,\bar{a})-f_{i,t}(s,\bar{a}))^{2}\right]\pi_{i,t}(0|s)+\textrm{Var}(\epsilon_{i,t})}{1-\pi_{i,t}(0|s)}$
$\qquad+\frac{\mathbb{E}_{p(s,\bar{a})}\left[(r_{i,t}(s,0)-f_{i,t}(s,0))^{2}\right][1-\pi_{i,t}(0|s)]+\textrm{Var}(\epsilon_{i,t})}{\pi_{i,t}(0|s)}$
@@
$\textrm{Var}\left(\tilde{R}_{i,t}^{f}(s,\bar{a})\right)$ $\leq\frac{o_{P}(k^{-1/2})\pi_{i,t}(0|s)+\textrm{Var}(\epsilon_{i,t})}{1-\pi_{i,t}(0|s)}+\frac{o_{P}(k^{-1/2})[1-\pi_{i,t}(0|s)]+\textrm{Var}(\epsilon_{i,t})+\textrm{Var}(\epsilon_{i,t})}{\pi_{i,t}(0|s)}$
@@
$\textrm{Var}\left(\tilde{R}_{i,t}^{f}(s,\bar{a})\right)$ $\leq\frac{Ck^{-1/2}\pi_{i,t}(0|s)+\textrm{Var}(\epsilon_{i,t})}{1-\pi_{i,t}(0|s)}+\frac{Ck^{-1/2}[1-\pi_{i,t}(0|s)]+\textrm{Var}(\epsilon_{i,t})}{\pi_{i,t}(0|s)}+2Ck^{-1/2}$
$\equiv v_{k}^{2}$
@@
$C$ $\equiv\max\left(A,16B^{2}K^{1/2}\right).$
Remark 1.
In the limit as $k\rightarrow\infty$
$\tilde{\sigma}^{2}_{i,t}\textrm{Var}\left(\tilde{R}_{i,t}^{f}(s,\bar{a})\right)$ $\leq\pi_{i,t}(0|s)\textrm{Var}(\epsilon_{i,t})+(1-\pi_{i,t}(0|s))\textrm{Var}(\epsilon_{i,t})$
$=\textrm{Var}(\epsilon_{i,t})$
@@
$v_{k}^{2}\equiv\frac{r_{i,t}(s,\bar{a})^{2}+\textrm{Var}(\epsilon_{i,t})}{1-\pi_{i,t}(0|s)}+\frac{r_{i,t}(s,0)^{2}+\textrm{Var}(\epsilon_{i,t})}{\pi_{i,t}(0|s)}-\Delta_{i,t}(s,\bar{a})^{2}$
@@
Here we collect some important results. We first adapt an important concentration inequality for regularized least-squares estimates. Our proof follows the same basic strategy as Abbasi-Yadkori et al. (2011), but with modifications due to (a) the use of weighted least squares (b) the use of pseudo-rewards to estimate differential rewards (c) replacing the scaled diagonal regularization with Laplacian regularization.
@@
Let $\hat{\theta}_{k}$ be the stage $k$ regularized least squares (RLS) estimate from Algorithm 1 and $\theta^{*}$ the ground truth. For any $\delta>0$, w.p. at least $1-\delta$ the estimates $\{\theta_{k}\}_{k=0}^{\infty}$ satisfies for any $\{x_{k}\}_{k=0}^{\infty}$,
$|x_{k}^{\top}(\hat{\theta}_{k}-\theta^{*})|$ $\leq\|x_{k}\|_{V_{k-1}^{-1}}\left(v_{k}\sqrt{2\log\left(\frac{\det(V_{k-1})^{1/2}\det(V_{0})^{-1/2}}{\delta}\right)}+\lambda_{\max}(V_{0})^{1/2}kB\right),$ (16)
where $v_{k}^{2}$ is the variance factor for the difference between the pseudo-reward and its mean at stage $k$. In particular, setting $x_{k}=V_{k-1}(\hat{\theta}_{k-1}-\theta^{*})$ implies
$\|\hat{\theta}_{k}-\theta^{*}\|_{V_{k-1}}$ $\leq v_{k}\sqrt{2\log\left(\frac{\det(V_{k-1})^{1/2}\det(V_{0})^{-1/2}}{\delta}\right)}+\lambda_{\max}(V_{0})^{1/2}kB$
@@
$=V_{k-1}^{-1}\xi_{k}+V_{k-1}^{-1}\Phi_{k}^{\top}W_{k}\Phi_{k}\theta^{*}$
$=V_{k-1}^{-1}\xi_{k}+V_{k-1}^{-1}(\Phi_{k}^{\top}W_{k}\Phi_{k}+V_{0})\theta^{*}-V_{k-1}^{-1}V_{0}\theta^{*}$
$=V_{k-1}^{-1}\xi_{k}+\theta^{*}-V_{k-1}^{-1}V_{0}\theta_{k}^{*}$
@@
$\hat{\theta}_{k}-\theta^{*}$ $=V_{k-1}^{-1}(\xi_{k}-V_{0}\theta^{*})$
@@
$|x_{k}^{\top}\hat{\theta}_{k}-x_{k}^{\top}\theta^{*}|$ $\leq\|x_{k}\|_{V_{k-1}^{-1}}(\|\xi_{k}\|_{V_{k-1}^{-1}}+\|V_{0}\theta*\|_{V_{k-1}^{-1}})$
@@
$\|V_{0}\theta^{*}\|_{V_{k-1}^{-1}}^{2}$ $=\theta_{k}^{*\top}V_{0}^{\top}V_{k-1}^{-1}V_{0}\theta^{*}$
$\leq\|V_{0}^{\top}V_{k-1}^{-1}V_{0}\|_{2}\|\theta^{*}\|_{2}^{2}$
$\leq\|V_{0}\|^{2}_{2}\|V_{k-1}^{-1}\|_{2}\|\theta^{*}\|_{2}^{2}$
$\leq\lambda_{\textrm{max}}(V_{0})\|\theta^{*}\|_{2}^{2}$
@@
$|x_{k}^{\top}\hat{\theta}_{k}-x_{k}^{\top}\theta^{*}|$ $\leq\|x_{k}\|_{V_{k-1}^{-1}}\left(v_{k}\sqrt{2\log\left(\frac{\textrm{det}(V_{k-1}^{-1})^{1/2}\textrm{det}(V_{0})^{-1/2}}{\delta}\right)}+\lambda_{\max}(V_{0})^{1/2}kB\right).$
Finally, setting $x_{k}=V_{k-1}(\hat{\theta}_{k-1}-\theta^{*})$ implies
$\|\hat{\theta}_{k-1}-\theta^{*}\|^{2}_{V_{k-1}}$ $=(\hat{\theta}_{k-1}-\theta^{*})^{\top}V_{k-1}^{\top}(\hat{\theta}_{k-1}-\theta^{*})$
$=|x_{k}^{\top}\hat{\theta}_{k}-x_{k}^{\top}\theta^{*}|$
$\leq\|x_{k}\|_{V_{k-1}^{-1}}\left(v_{k}\sqrt{2\log\left(\frac{\textrm{det}(V_{k-1}^{-1})^{1/2}\textrm{det}(V_{0})^{-1/2}}{\delta}\right)}+\lambda_{\max}(V_{0})^{1/2}kB\right)$
$=\|V_{k-1}(\hat{\theta}_{k-1}-\theta^{*})\|_{V_{k-1}^{-1}}\left(v_{k}\sqrt{2\log\left(\frac{\textrm{det}(V_{k-1}^{-1})^{1/2}\textrm{det}(V_{0})^{-1/2}}{\delta}\right)}+\lambda_{\max}(V_{0})^{1/2}kB\right)$
$=\|\hat{\theta}_{k-1}-\theta^{*}\|_{V_{k-1}}\left(v_{k}\sqrt{2\log\left(\frac{\textrm{det}(V_{k-1}^{-1})^{1/2}\textrm{det}(V_{0})^{-1/2}}{\delta}\right)}+\lambda_{\max}(V_{0})^{1/2}kB\right)$
and dividing both sides by $\|\hat{\theta}_{k-1}-\theta^{*}\|_{V_{k-1}}$ implies
$\|\hat{\theta}_{k}-\theta^{*}\|_{V_{k-1}}$ $\leq v_{k}\sqrt{2\log\left(\frac{\det(V_{k-1})^{1/2}\det(V_{0})^{-1/2}}{\delta}\right)}+\lambda_{\max}(V_{0})^{1/2}kB$
@@
$|x_{k}^{\top}(\hat{\theta}_{k}-\theta^{*})|$ $\leq\|x_{k}\|_{V_{k-1}^{-1}}\left(\left(\frac{C}{k^{1/2}}+\sigma^{2}c^{2}\right)\sqrt{2\log\left(\frac{\det(V_{k-1})^{1/2}\det(V_{0})^{-1/2}}{\delta/2}\right)}+\lambda_{\max}(V_{0})^{1/2}kB\right),$ (17)
In particular, setting $x_{k}=V_{k-1}(\hat{\theta}_{k-1}-\theta^{*})$ implies
$\|\hat{\theta}_{k}-\theta^{*}\|_{V_{k-1}}$ $\leq\left(\frac{C}{k^{1/2}}+\sigma^{2}c^{2}\right)\sqrt{2\log\left(\frac{\det(V_{k-1})^{1/2}\det(V_{0})^{-1/2}}{\delta/2}\right)}+\lambda_{\max}(V_{0})^{1/2}kB$
@@
Let $\lambda\geq 1$ and $\gamma\geq 1$. For any arbitrary sequence $(x_{i,t})_{(i,t)\in\mathcal{O}_{k}}$, let
$V_{k+1}\equiv\sum_{(i,t)\in\mathcal{O}_{k}}\tilde{\sigma}_{i,t}^{2}\phi(x_{i,t})\phi(x_{i,t})^{\top}+V_{0},$
@@
$\sum_{k=1}^{K}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\|\phi(x_{i,t})\|_{V_{k}^{-1}}^{2}\leq 2c\log\left(\frac{\det(V_{K+1})}{\det(V_{0})}\right).$
@@
$\sum_{k=1}^{K}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\tilde{\sigma}_{i,t}^{2}\|\phi(x_{i,t})\|_{V_{k}^{-1}}^{2}\leq 2\log\left(\frac{\det(V_{K+1})}{\det(V_{0})}\right).$
@@
$\sum_{k=1}^{K}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\|\phi(x_{i,t})\|_{V_{k}^{-1}}^{2}$ $\leq 2c\log\left(\frac{\det(V_{K+1})}{\det(V_{0})}\right)$
as desired. Now,
$tr(\tilde{\sigma}_{i,t}^{2}\phi(x_{i,t})\phi(x_{i,t})^{\top})$ $=\tilde{\sigma}_{i,t}^{2}tr(\phi(x_{i,t})^{\top}\phi(x_{i,t}))$
$=\tilde{\sigma}_{i,t}^{2}\|\phi(x_{i,t})^{\top}\|^{2}$
$=3\sigma_{i,t}^{2}\|x_{i,t}\|^{2}$
$\leq\frac{3}{4}$
and $tr(L_{\otimes})=tr(L)\cdot tr(I)=KMd$. Then by Hadamard’s inequality we have
@@
$=2Kd\log\left(\frac{1}{2Kd}\left[\gamma 2Kd+\lambda 2KMd+\frac{3}{4}K\cdot(K+1)/2\right]\right)$
$=2Kd\log\left(\left[\gamma+\lambda M+3\frac{K+1}{16d}\right]\right)$